If a car,with an initial velocity of 20m/s, accelerates at a rate of 5m/s2 for 3 seconds, what will its final velocity be?


1
Expert's answer
2021-01-31T03:43:08-0500

Initial velocity of car, "u=20m\/s"

Acceleration of car, "a=5m\/s^2"

Time, "t=3s"

Velocity after 3s, "v=u+at=20+(5*3)=35m\/s"